# OLMo[Uncaptioned image] : Accelerating the Science of Language Models

### 01 Feb 2024 v1

@@ 1 Introduction
OLMo releases the whole framework from data to training to evaluation tools: multiple training checkpoints across multiple hardware types, training logs, and exact datasets used, with a permissive license. We are not the only team to do this; recent work from LLM360 targets similar goals (Liu et al., 2023). OLMo narrows the gap from their models to state-of-the-art capabilities of models like LLaMA2. This project has benefited from lessons learned from all of these previous efforts with their varying degrees of openness, and we believe that a large, diverse population of open models is the best hope for scientific progress on understanding language models and engineering progress on improving their utility.
@@ 2.3 Evaluation
We perform model evaluation at two stages: online evaluation to make decisions for model design and offline evaluation to evaluate model checkpoints. For offline evaluation, we use the Catwalk framework (Groeneveld et al., 2023), our publicly available evaluation tool with access to a wide range of datasets and task formats. Using Catwalk, we perform downstream evaluation as well as intrinsic language modeling evaluation on our new perplexity benchmark, Paloma (Magnusson et al., 2023).
For both downstream and perplexity evaluation, we use our fixed evaluation pipeline to compare results against several publicly available models.
@@
Following much previous work (Brown et al., 2020; Black et al., 2022; Touvron et al., 2023a, b, inter alia), we report zero-shot performance on a set of downstream tasks. Our evaluation suite consists of 9 core tasks corresponding closely to the commonsense reasoning task set reported by Touvron et al. (2023a) and Touvron et al. (2023b) (see Table 6 for a list of tasks). Given the scale of the models being evaluated, such tasks were selected at the beginning of model development due to their naturalness (e.g., all can formulated as text completion scoring tasks) and ability to provide meaningful signals throughout training (see Figure 1).
@@ 4 Results
The checkpoint used for evaluating OLMo-7B is trained until 2.46T tokens on the Dolma (Soldaini et al., 2024) dataset with a linear learning rate decay schedule mentioned in Section 3.2. In our experiments, we find that tuning this checkpoint further on Dolma dataset for 1000 steps with the learning rate linearly decayed to 0 boosts model performance on perplexity and end-task evaluation suites described in Section 2.3. We compare OLMo with other publicly available models including LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), MPT-7B (MosaicML NLP Team, 2023), Pythia-6.9B (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023).
@@ 4.1 Downstream evaluation
Our core downstream evaluation suite (see Table 6) consists of: arc (both arc_easy and arc_challenge) (Clark et al., 2018), boolq (Clark et al., 2019), openbookqa (Mihaylov et al., 2018), sciq (Welbl et al., 2017), hellaswag (Zellers et al., 2019), piqa (Bisk et al., 2020), copa (Roemmele et al., 2011) and winogrande (Sakaguchi et al., 2021). In Appendix A, we also report results on an additional set of auxiliary tasks outside of our core evaluation set that we found to have less stable performance trends (see Figure 4). We note that our downstream evaluation suite is still under development and that additional results and analysis will be reported in a future version.
In all cases, we perform zero-shot evaluation using the rank classification approach popularized by Brown et al. (2020). Under this approach, candidate text completions (e.g., different multiple-choice options) are ranked by likelihood (usually normalized by some normalization factor), and prediction accuracy is reported. While Catwalk implements several common likelihood normalization strategies, including normalizing by number of tokens (per-token normalization) (Brown et al., 2020; Liang et al., 2022), by number of characters (per-character normalization) (Gao et al., 2023), as well as incorporating an answer’s unconditional likelihood (Brown et al., 2020), we selected the normalization strategies for each dataset separately. Specifically, we used unconditional normalization for arc and openbookqa, per-token normalization for hellaswag, piqa, and winogrande and no normalization for boolq, copa, and sciq (i.e., tasks formulated as single token prediction tasks).
7B Models arc challenge arc easy boolq copa hella- swag open bookqa piqa sciq wino- grande avg. Falcon 47.5 70.4 74.6 86.0 75.9 53.0 78.5 93.9 68.9 72.1 LLaMA 44.5 57.0 73.1 85.0 74.5 49.8 76.3 89.5 68.2 68.7 LLaMA2 39.8 57.7 73.5 87.0 74.5 48.4 76.4 90.8 67.3 68.4 MPT 46.5 70.5 74.2 85.0 77.6 48.6 77.3 93.7 69.9 71.5 Pythia 44.2 61.9 61.1 84.0 63.8 45.0 75.1 91.1 62.0 65.4 RPJ-INCITE 42.8 68.4 68.6 88.0 70.3 49.4 76.0 92.9 64.7 69.0 [] OLMo-7B 48.5 65.4 73.4 90.0 76.4 50.4 78.4 93.8 67.9 71.6
Table 6: Zero-shot evaluation of OLMo-7B and 6 other publicly available comparable model checkpoints on 9 core tasks from the downstream evaluation suite described in Section 2.3. For OLMo-7B, we report results for the 2.46T token checkpoint.
@@ Results
Table 6 summarizes the result of zero-shot evaluation of OLMo-7B and compares it against 6 other publicly available models of comparable size. We report results on 9 core tasks from our evaluation suite described in Section 2.3. Our OLMo-7B checkpoint outperforms all other publicly available models on 2 end-tasks and remains in top-3 on 8/9 end-tasks from the evaluation suite. On aggregate, OLMo-7B is competitive against all 6 publicly available model checkpoints in our comparison table.
In Figure 1 we plot the accuracy score progression of 9 core end-tasks. All tasks, except OBQA, show an upward trend in accuracy numbers as OLMo-7B is trained on more tokens. A sharp upward tick in accuracy of many tasks between the last and the second to last step shows us the benefit of linearly reducing the LR to 0 over the final 1000 training steps. See Table 8 in Appendix A for additional evaluation results and discussion.
Figure 1: Accuracy score progression of OLMo-7B on 9 core end-tasks score from Catwalk evaluation suite described in Section 2.3. We can see the benefit of decaying LR to 0 in the final 1000 steps of training on 7/9 end-tasks.
@@ 5 Artifacts Released
    The training and modeling code. [^12] https://github.com/allenai/OLMo.
@@
    The trained model weights for the 7B model [^13] https://huggingface.co/allenai/OLMo-7B, 7B-twin-2T [^14] https://huggingface.co/allenai/OLMo-7B-Twin-2T, and the 1B model [^15] https://huggingface.co/allenai/OLMo-1B. For all the models, we release not only the final model weights but also 500+ intermediate checkpoints at intervals of 1000 steps.
@@
  4. 4.
    Dolma’s toolkit to construct new datasets  [^17] https://github.com/allenai/dolma, and WIMBD (Elazar et al., 2023) for dataset analysis [^18] https://github.com/allenai/wimbd.
@@
We intend to follow up on this release with another one soon that includes the following:
  1. 1.
    Training logs, ablations, and findings.
  2. 2.
    Weights & Biases logs for our training runs.
  3. 3.
    Adapted OLMo with instruction-tuning and RLHF, including its training and evaluation code and data using our Open Instruct [^22] https://github.com/allenai/open-instruct library (Wang et al., 2023; Ivison et al., 2023).
@@ 6 License
Our goal is to facilitate scientific development and empower the scientific community, so we favor permissive licenses that give users flexibility in using our resources and artifacts. As such, all code and weights are released under the Apache 2.0 License. [^23] http://www.apache.org/licenses/LICENSE-2.0 Some licenses used by other organizations for recent model releases prohibit using the outputs from their models to train artificial intelligence or machine learning systems, while we expressly allow users to do so. We also do not limit commercial use. We hope that our models can make other models better. We recognize that the risk for misuse of our models is relatively low, as language models that have not been adapted as chatbots have primarily been used as scientific artifacts not as products with broad public adoption (our models have not been adapted as chatbots). In addition, over the past year there have been a number of comparable models released with very permissive licenses, so using a more strict license for our work will not remove the overall risk in the field. We believe this tradeoff on the side of being more open is the best option.
@@ 7 Conclusion and Future Work
Contributors to evaluation suite and tooling include Akshita Bhagia, Arman Cohan, Pradeep Dasigi, Jesse Dodge, Dirk Groeneveld, Yuling Gu, Tushar Khot, Kyle Richardson, Oyvind Tajford, and Pete Walsh.
@@ Acknowledgements
OLMo would not have been possible without the support of many individuals and institutions. The experimental components of this work were made possible through a partnership with AMD and CSC, enabling use of the LUMI supercomputer, and Kempner Institute at Harvard University. We thank Jonathan Frankle and the team at MosaicML (now Databricks) for sharing their experiences with FSDP, and building the code base that OLMo is based on. We thank our teammates Taira Anderson, Michelle Benedict, Jon Borchardt, Evie Cheng, Arnavi Chheda, Johann Dahm, Matt Latzke, Kelsey MacMillan, Aaron Sarnat, Carissa Schoenick, Sam Skjonsberg, Michael Schmitz, Michael Wilson, Caitlin Wittlif, and the entire IT team, for their help with the website, design, internal and external communications, budgeting, and other activities that supported smooth progress on this project. Finally, we also express gratitude for the helpful discussions and feedback from our teammates at AI2 and close collaborators, including Prithviraj (Raj) Ammanabrolu, Peter Clark, Nicole DeCario, Doug Downey, Ali Farhadi, Ian Ferreira, Väinö Hatanpää, Sham M. Kakade, Julien Launay, Sydney Levine, Pekka Manninen, Franzi Roessner, Maarten Sap, Ludwig Schmidt, and Yulia Tsvetkov.
@@ Appendix A Additional Evaluation
Next, in Table 8, we provide results from zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 9 in our core evaluation suite. These tasks are headqa_en [Vilares and Gómez-Rodríguez, 2019], logiqa [Liu et al., 2020], mrpc [Dolan and Brockett, 2005], qnli [Wang et al., 2018], wic [Pilehvar and Camacho-Collados, 2018], and wnli [Wang et al., 2018].
headqa_en logiqa mrpc qnli wic wnli avg. Falcon-7B 38.6 23.7 62.8 49.8 49.5 47.9 45.4 LLaMA-7B 37.0 20.9 68.4 48.4 50.0 56.3 46.8 LLaMA2-7B 38.3 22.7 64.2 49.2 50.3 49.3 45.7 MPT-7B 37.4 22.9 67.7 52.1 48.1 47.9 46.0 Pythia-6.9B 40.1 21.5 65.4 53.8 55.0 38.0 45.6 RPJ-INCITE-7B 36.9 27.8 58.8 53.8 48.9 57.8 47.3 [] OLMo-7B 37.3 23.4 68.4 49.1 50.2 56.3 47.5
Table 8: Zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 9 present in our core evaluation suite. Once again, we compare OLMo-7B to 6 other model checkpoints which are publicly available. We find that OLMo-7B outperforms the other models on aggregate taken over 6 additional end-tasks from this table, however these tasks were also found to provide limited signal during training (see Figure 4).

### 28 Feb 2024 v3

@@ Abstract
[Uncaptioned image] W&B Logs https://wandb.ai/ai2-llm/OLMo-7B/reports/OLMo-7B--Vmlldzo2NzQyMzk5
@@ 1 Introduction
With OLMo, we release the whole framework from data to training to evaluation tools: multiple training checkpoints across multiple hardware types, training logs, and exact datasets used, with a permissive license. We are not the only team to do this; recent work from LLM360 targets similar goals (Liu et al., 2023). OLMo narrows the gap from their models to state-of-the-art capabilities of models like LLaMA2. This project has benefited from lessons learned from all of these previous efforts with their varying degrees of openness, and we believe that a large, diverse population of open models is the best hope for scientific progress on understanding language models and engineering progress on improving their utility.
@@ 2.2 Pretraining Data: Dolma
2.3 Adaptation
Pretrained models are not always used as-is, but rather further fine-tuned to improve their performance, safety, and usability. Often models are first trained to follow instructions (Mishra et al., 2022; Wei et al., 2022; Sanh et al., 2022), and then further trained on human preferences (Ouyang et al., 2022) to improve the quality of their generations. We showcase the efficacy of using OLMo as a base model for further fine-tuning by training OLMo to be a general chat assistant following our Open Instruct (Tülu) data and training setup (Ivison et al., 2023). Our approach involves first performing instruction fine-tuning with a mixture of distilled and human-written instruction data and then further aligning the model with distilled preference data using Direct Preference Optimization (DPO) (Rafailov et al., 2023). We experimented with mixing the Tulu instruction data at the end of pretraining, as done in recent models such as DeepSeek-AI et al. (2024), but did not have conclusive findings.
@@ 2.4 Evaluation
We perform base model evaluation at two stages: online evaluation to make decisions for model design and offline evaluation to evaluate model checkpoints. For the offline stage, we use the Catwalk framework (Groeneveld et al., 2023), a publicly available evaluation tool with access to a wide range of datasets and task formats. Using Catwalk, we perform downstream evaluation as well as intrinsic language modeling evaluation on the new perplexity benchmark, Paloma (Magnusson et al., 2023).
For both downstream and perplexity evaluation, we use our fixed evaluation pipeline to compare results against publicly available models. We also report a separate evaluation of our adapted model.
@@
Following much previous work (Brown et al., 2020; Black et al., 2022; Touvron et al., 2023a, b, inter alia), we report zero-shot performance on a set of downstream tasks. Our evaluation suite consists of 8 core tasks corresponding closely to the commonsense reasoning task set reported by Touvron et al. (2023a) and Touvron et al. (2023b) (see Table 6 for a list of tasks). Given the scale of the models being evaluated, such tasks were selected at the beginning of model development due to their naturalness (e.g., all can formulated as text completion scoring tasks) and ability to provide meaningful signals throughout training (see Figure 1).
@@
Adaptation Evaluation
We also follow our Open Instruct evaluation suite Wang et al. (2023); Ivison et al. (2023) to evaluate OLMo after instruction fine-tuning and DPO training using our We focus on evaluations around model chat capabilities and safety to showcase the efficacy of using OLMo as a base for further fine-tuning.
@@ 4 Results
The checkpoint used for evaluating OLMo-7B is trained until 2.46T tokens on the Dolma (Soldaini et al., 2024) dataset with a linear learning rate decay schedule mentioned in Section 3.2. In our experiments, we find that tuning this checkpoint further on the Dolma dataset for 1000 steps with the learning rate linearly decayed to 0 boosts model performance on perplexity and end-task evaluation suites described in Section 2.4. We compare OLMo with other publicly available models including LLaMA-7B (Touvron et al., 2023a), LLaMA2-7B (Touvron et al., 2023b), MPT-7B (MosaicML NLP Team, 2023), Pythia-6.9B (Biderman et al., 2023), Falcon-7B (Almazrouei et al., 2023) and RPJ-INCITE-7B (Together Computer, 2023).
@@ 4.1 Downstream evaluation
Our core downstream evaluation suite (see Table 6) consists of: arc (both arc_easy and arc_challenge) (Clark et al., 2018), boolq (Clark et al., 2019), openbookqa (Mihaylov et al., 2018), sciq (Welbl et al., 2017), hellaswag (Zellers et al., 2019), piqa (Bisk et al., 2020), and winogrande (Sakaguchi et al., 2021). In Appendix A, we also report results on an additional set of auxiliary tasks outside of our core evaluation set that we found to have less stable performance trends (see Figure 4).
In all cases, we perform zero-shot evaluation using the rank classification approach popularized by Brown et al. (2020). Under this approach, candidate text completions (e.g., different multiple-choice options) are ranked by likelihood (usually normalized by some normalization factor), and prediction accuracy is reported. While Catwalk implements several common likelihood normalization strategies, including normalizing by number of tokens (per-token normalization) (Brown et al., 2020; Liang et al., 2022), by number of characters (per-character normalization) (Gao et al., 2023), as well as incorporating an answer’s unconditional likelihood (Brown et al., 2020), we selected the normalization strategies for each dataset separately. Specifically, we used unconditional normalization for arc and openbookqa, per-token normalization for hellaswag, piqa, and winogrande and no normalization for boolq, and sciq (i.e., tasks formulated as single token prediction tasks).
7B Models
arc
challenge
arc
easy
boolq
hella-
swag
open
bookqa
piqa sciq
wino-
grande
avg.
Falcon 47.5 70.4 74.6 75.9 53.0 78.5 93.9 68.9 70.3
LLaMA 44.5 67.9 75.4 76.2 51.2 77.2 93.9 70.5 69.6
Llama 2 48.5 69.5 80.2 76.8 48.4 76.7 94.5 69.4 70.5
MPT 46.5 70.5 74.2 77.6 48.6 77.3 93.7 69.9 69.8
Pythia 44.1 61.9 61.1 63.8 45.0 75.1 91.1 62.0 63.0
RPJ-INCITE 42.8 68.4 68.6 70.3 49.4 76.0 92.9 64.7 66.6
[] OLMo-7B 48.5 65.4 73.4 76.4 50.4 78.4 93.8 67.9 69.3
Table 6: Zero-shot evaluation of OLMo-7B and 6 other publicly available comparable model checkpoints on 8 core tasks from the downstream evaluation suite described in Section 2.4. For OLMo-7B, we report results for the 2.46T token checkpoint.
@@ Results
Table 6 summarizes the result of zero-shot evaluation of OLMo-7B and compares it against 6 other publicly available models of comparable size. We report results on 8 core tasks from our evaluation suite described in Section 2.4. On aggregate, OLMo-7B is competitive against all 6 publicly available model checkpoints in our comparison table.
In Figure 1 we plot the accuracy score progression of 8 core end-tasks. All tasks, except OBQA, show an upward trend in accuracy numbers as OLMo-7B is trained on more tokens. A sharp upward tick in accuracy of many tasks between the last and the second to last step shows us the benefit of linearly reducing the LR to 0 over the final 1000 training steps. See Table 9 in Appendix A for additional evaluation results and discussion.
Figure 1: Accuracy score progression of OLMo-7B on 8 core end-tasks score from Catwalk evaluation suite described in Section 2.4. We can see the benefit of decaying LR to 0 in the final 1000 steps of training on most tasks.
@@
4.3 Adaptation Evaluation
Model MMLU AlpacaEval ToxiGen TruthfulQA 0-shot $\uparrow$ %win $\uparrow$ % Toxic $\downarrow$ %Info+True $\uparrow$ OLMo (base) 28.3 - 81.4 31.6 MPT Chat 33.8 46.8 0.1 42.7 Falcon Instruct 25.2 14.0 70.7 27.2 RPJ-INCITE Chat 27.0 38.0 46.4 53.0 Llama-2-Chat 46.8 87.3 0.0 26.3 Tülu 2 50.4 73.9 7.0 51.7 Tülu 2+DPO 50.7 85.1 0.5 - * [] OLMo +SFT 47.3 57.0 14.4 41.2 [] OLMo +SFT+DPO 46.2 69.3 1.7 52.0
Table 7: Evaluation of various instruction-tuned 7B models, including OLMo-7B and before and after adaptation training. Lower is better for ToxiGen and higher is better for other metrics. We provide a detailed description of models and metrics in Appendix. C. * Following Ivison et al. (2023), we do not report Tülu 2 TruthfulQA scores due to test set contamination.
Setup
We evaluate OLMo before adaptation, and after both the supervised fine-tuning and DPO training stage, focusing on the safety and chat evaluations used by Wang et al. (2023). We additionally compare to officially released instruction-tuned variants of the models from Table 6. We finally also compare to Tülu 2 models to compare against models trained using the same post-training data mixes and procedures.
Results
We find that instruction tuning considerably improves the performance and safety of OLMo, increasing MMLU performance by a wide margin and improving ToxiGen and TruthfulQA scores - especially after DPO training. Additionally, we find that OLMo outperforms most other chat variants after both initial instruction tuning (OLMo +SFT) and additional preference alignment (OLMo +SFT+DPO), highlighting both the strength of OLMo as a base model and the strength of the Tülu mix used to perform adaptation training. However, we find there is still a gap with Tülu 2, which is trained by applying the Tülu mix on Llama 2. This gap may be due to test set contamination in Llama 2777Touvron et al. (2023b) report that Llama 2 was pretrained on data contaminated with MMLU test data. and because the Tülu mix was primarily designed for Llama models - we will investigate the cause of this gap in future work. Overall, we see that OLMo greatly benefits from additional tuning and serves as a strong base model for downstream applications.
@@ 5 Artifacts Released
    The training and modeling code. [^13] https://github.com/allenai/OLMo
@@
    The trained model weights for the 7B model, [^14] https://huggingface.co/allenai/OLMo-7B 7B-twin-2T, [^15] https://huggingface.co/allenai/OLMo-7B-Twin-2T and the 1B model. [^16] https://huggingface.co/allenai/OLMo-1B For all the models, we release not only the final model weights but also 500+ intermediate checkpoints at intervals of 1000 steps.
@@
    Adapted OLMo-7B with instruction-tuning, 7B-SFT [^17] https://huggingface.co/allenai/OLMo-7B-SFT, and RLHF, 7B-Instruct [^18] https://huggingface.co/allenai/OLMo-7B-Instruct including its training and evaluation code and data using our Open Instruct [^19] https://github.com/allenai/open-instruct library (Wang et al., 2023; Ivison et al., 2023).
  4. 4.
@@
    Dolma’s toolkit to construct new datasets, [^21] https://github.com/allenai/dolma and WIMBD (Elazar et al., 2023) for dataset analysis. [^22] https://github.com/allenai/wimbd
  6. 6.
@@
  7. 7.
    The complete set of metrics logged to Weights & Biases during training. [^26] https://wandb.ai/ai2-llm/OLMo-7B/reports/OLMo-7B--Vmlldzo2NzQyMzk5
We intend to follow up on this release with further training logs, ablations, and findings.
@@ 6 License
Our goal is to facilitate scientific development and empower the scientific community, so we favor permissive licenses that give users flexibility in using our resources and artifacts. As such, all code and weights are released under the Apache 2.0 License. [^27] http://www.apache.org/licenses/LICENSE-2.0 Some licenses used by other organizations for recent model releases prohibit using the outputs from their models to train artificial intelligence or machine learning systems, while we expressly allow users to do so. We also do not limit commercial use. We hope that our models can make other models better. We recognize that the risk for misuse of our models is relatively low since they are mainly designed as scientific artifacts not as products with broad public adoption (our models have not been adapted as chatbots). In addition, over the past year there have been a number of comparable models released with very permissive licenses, so using a more strict license for our work will not remove the overall risk in the field. We believe this tradeoff on the side of being more open is the best option.
@@ 7 Conclusion and Future Work
Contributors to evaluation suite and tooling include Akshita Bhagia, Arman Cohan, Pradeep Dasigi, Jesse Dodge, Dirk Groeneveld, Yuling Gu, Tushar Khot, Ian Magnusson, Kyle Richardson, Oyvind Tajford, and Pete Walsh.
@@ Acknowledgements
OLMo would not have been possible without the support of many individuals and institutions. The experimental components of this work were made possible through a partnership with AMD and CSC, enabling use of the LUMI supercomputer, and Kempner Institute at Harvard University. We thank Jonathan Frankle and the team at MosaicML (now Databricks) for sharing their experiences with FSDP, and building the code base that OLMo is based on. We thank our teammates Taira Anderson, Michelle Benedict, Jon Borchardt, Evie Cheng, Arnavi Chheda, Johann Dahm, Matt Latzke, Kelsey MacMillan, Aaron Sarnat, Carissa Schoenick, Sam Skjonsberg, Michael Schmitz, Michael Wilson, Caitlin Wittlif, and the entire IT team, for their help with the website, design, internal and external communications, budgeting, and other activities that supported smooth progress on this project. Finally, we also express gratitude for the helpful discussions and feedback from our teammates at AI2 and close collaborators, including Prithviraj (Raj) Ammanabrolu, Peter Clark, Nicole DeCario, Doug Downey, Ali Farhadi, Ian Ferreira, Väinö Hatanpää, Sham M. Kakade, Julien Launay, Sydney Levine, Pekka Manninen, Franzi Roessner, Maarten Sap, Ludwig Schmidt, Yulia Tsvetkov, and Daniel S. Weld.
@@ Appendix A Additional Evaluation
Next, in Table 9, we provide results from zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 8 in our core evaluation suite. These tasks are headqa_en [Vilares and Gómez-Rodríguez, 2019], logiqa [Liu et al., 2020], mrpc [Dolan and Brockett, 2005], qnli [Wang et al., 2018], wic [Pilehvar and Camacho-Collados, 2018], and wnli [Wang et al., 2018].
headqa_en logiqa mrpc qnli wic wnli avg. Falcon-7B 38.6 23.7 62.8 49.8 49.5 47.9 45.4 LLaMA-7B 38.7 19.5 68.6 50.1 49.1 52.1 46.4 LLaMA2-7B 39.5 26.1 69.1 49.4 49.8 45.1 46.5 MPT-7B 37.4 22.9 67.7 52.1 48.1 47.9 46.0 Pythia-6.9B 40.1 21.5 65.4 53.8 55.0 38.0 45.6 RPJ-INCITE-7B 36.9 27.8 58.8 53.8 48.9 57.8 47.3 [] OLMo-7B 37.3 23.4 68.4 49.1 50.2 56.3 47.5
Table 9: Zero-shot evaluation of OLMo-7B on 6 additional end-tasks apart from the 8 present in our core evaluation suite. Once again, we compare OLMo-7B to 6 other model checkpoints which are publicly available. We find that OLMo-7B outperforms the other models on aggregate taken over 6 additional end-tasks from this table, however these tasks were also found to provide limited signal during training (see Figure 4).
@@
Appendix B Adaptation Training Details
We use the following hyperparameters when instruction tuning OLMo. These were chosen through small pilot experiments.
  • Learning Rate: $2\times 10^{-6}$
  • Epochs: 3
  • Warmup: Linear warmup for the first 3% of total training time, and then linear cooldown to a learning rate of 0 over the remaining steps.
  • Weight Decay: 0
  • Gradient clipping: 0
  • Maximum sequence length: 2048
After instruction finetuning, we then use the following hyperparameters for DPO training, following Ivison et al. [2023]:
  • Learning Rate: $5\times 10^{-7}$
  • $\beta$: 0.1
  • Epochs: 3
  • Warmup: Linear warmup for the first 10% of total training time, and then linear cooldown to a learning rate of 0 over the remaining steps.
  • Weight Decay: 0
  • Gradient clipping: 0
  • Maximum sequence length: 2048
Appendix C Adaptation Evaluation and Model details
We choose the models in Table 7 by choosing the ‘canonical’ best versions (that is, the best instruction-tuned or otherwise adapted models released by the same organisation) of the base models we compare against in Table 6. We additionally compare to Tülu 2 to show the current best models trained using the Tülu mix used to finetune OLMo. We display evaluations on MMLU, AlpacaEval, ToxiGen, and Truthfulness to focus on displaying how instruction tuning can generally help capabilities (MMLU), how the models perform in an open-ended chat setting (AlpacaEval), and to test how instruction tuning aids in model safety and truthfulness (AlpacaEval, ToxiGen). We additionally report OLMo’s performance over the entire Tülu evaluation suite in Table 10.
Model MMLU GSM8k BBH TydiQA Codex-Eval AlpacaEval ToxiGen TruthfulQA 0-shot 8-shot CoT 3-shot CoT 1-shot Pass@10 %win % Toxic % Info + True OLMo-7B 28.3 8.5 31.7 32.3 21.4 - 81.4 31.6 +SFT 47.3 15.5 36.9 35.2 28.6 57.0 14.4 41.2 +SFT+DPO 46.1 11.0 35.8 21.7 27.8 69.3 1.7 52.0
Table 10: Evaluation of OLMo-7B models before and after instruction finetuning and DPO training on the full Tülu evaluation suite. Lower is better for ToxiGen and higher is better for other metrics.
We provide a brief description of each model evaluated in Table 7 below. For all models, we use the provided chat template for prompt formatting when available.
  • MPT Chat: A version of MPT 7B finetuned on the ShareGPT-Vicuna [Chiang et al., 2023], HC3 [Guo et al., 2023], Alpaca [Taori et al., 2023], HH-RLHF [Bai et al., 2022], and Evol-Instruct [Xu et al., 2024] datasets. Retrieved from https://huggingface.co/mosaicml/mpt-7b-chat.
  • Falcon Instruct: A version of Falcon 7B finetuned on the Baize [Xu et al., 2023], GPT4All [Anand et al., 2023], GPTeacher [Teknium1, 2023], and Refined-Web English [Penedo et al., 2023] datasets. Retrieved from https://huggingface.co/tiiuae/falcon-7b-instruct.
  • RPJ-INCITE Chat: A version of RPJ-INCITE 7B finetuned on the OASST1 [Köpf et al., 2023] and Dolly V2 [Conover et al., 2023] datasets. Retrieved from https://huggingface.co/togethercomputer/RedPajama-INCITE-7B-Chat.
  • Llama-2 Chat: A version of Llama 2 7B finetuned on a mixture of instruction datasets and further trained with RLHF. We refer the reader to Touvron et al. [2023b] for further details.
  • Tülu 2: A version of Llama 2 7B finetuned on a mixture of instruction datasets (the Tülu 2 mix). We refer the reader to Ivison et al. [2023] for further details.
  • Tülu 2+DPO: Tülu 2 further trained with DPO on the UltraFeedback dataset [Cui et al., 2023]. We refer the reader to Ivison et al. [2023] for further details.
  • OLMo +SFT: A version of OLMo 7B fintuned on the same data as Tülu 2.
  • OLMo +SFT+DPO: OLMo +SFT further trained with DPO on the UltraFeedback dataset [Cui et al., 2023].
We additionally provide a brief description of each evaluation setting from Table 7:
  • MMLU: We use the official MMLU [Hendrycks et al., 2021] evaluation script and prompts available at https://github.com/hendrycks/test, with modifications to allow for batch processing. We evaluate using 0 few-shot examples, following the original setup of MMLU. We report average accuracy across test examples.
  • ToxiGen: We follow the setup in Touvron et al. [2023b], but use the original set of prompts from Hartvigsen et al. [2022], which are designed to elicit toxic generations for certain groups. We take only the prompts designed to produce toxic language (‘hateful’ prompts) and use 500 prompts per group to reduce evaluation costs. For base language models, we pass in the original ToxiGen prompts unchanged and greedily decode up to the first new line (or a maximum of 512 tokens). For instruction-tuned models, we place the prompt in the corresponding template, and ask the model to complete the prompt, until the model generates a stop token (or a maximum of 512 tokens). We pass the generated text into a roberta-large model trained to detect toxic content finetuned as part of Hartvigsen et al. [2022] [^28] https://huggingface.co/tomh/toxigen_roberta. We then report the percentage of generations deemed toxic by the classifier.
  • TruthfulQA: Following Touvron et al. [2023b], we mainly use the generation setting of TruthfulQA [Lin et al., 2022]. The TruthfulQA dataset contains 818 questions, which are used to prompt the tested model to generate answers. We use the default QA prompt format with 6 in-context QA examples. We follow the official script in their official implemention [^29] https://github.com/sylinrl/TruthfulQA/ to do greedy decoding and answer postprocessing. We train two LLaMA 2-based classifiers for judging the truthfulness and informativeness of the model response, due to the deprecation of GPT-3 making exact replication of the original TruthfulQA evaluation infeasible. We find that the LLaMA 2 judges are generally able to match the performance of the original GPT-3-based judges used by Lin et al. [2022]. We report the rate of the responses being truthful and informative (Informative and Truthful) following Touvron et al. [2023b]. We only report the % Informative and Truthful as our primary metric.
  • AlpacaEval: We use the package provided by Li et al. [2023], following the default setup which asks the evaluated model to generate responses for 805 prompts and employ GPT-4 to compare the response with Davinci-003. We employ the “alpaca_eval_gpt4” annotator. We allow the evaluated model to generate up to 2048 tokens, without specifying special stop sequences. The reported win-rate is the percentage of model generations that GPT-4 reports as being preferred over the generations from Davinci-003.